# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ Abstract
There exist two main approaches for non-intrusive implementations of approximate balanced truncation within the Loewner framework: the quadrature-based method [1] and the Alternating Direction Implicit (ADI)-based method [2]. Both approaches rely solely on samples of the transfer function to construct truncated balanced models, eliminating the need for access to the original model’s state-space realization. Recently, the quadrature-based approach has been extended to various generalizations of balanced truncation, including positive-real balanced truncation, bounded-real balanced truncation, and balanced stochastic truncation. While this extension [3] is theoretically non-intrusive—meaning it does not require the original state-space realization—it depends on samples of spectral factorizations of the transfer function. Since practical methods for obtaining such samples are currently unavailable, this extension remains largely a theoretical contribution.
In this work, we present a non-intrusive ADI-type framework for these generalized balanced truncation methods that requires only samples of the original transfer function for implementation. To achieve this, we first demonstrate how the approximations of ADI-type method for Riccati equations (RADI) [4] can be equivalently obtained via projection-based interpolation. Leveraging this insight, we develop several non-intrusive, data-driven algorithms that not only satisfy the desired interpolation conditions but also preserve key system properties such as positive-realness, bounded-realness, and minimum phase. Notably, these algorithms rely solely on solving Lyapunov and Sylvester equations, avoiding computationally intensive Riccati equations or linear matrix inequalities. Additionally, we propose non-intrusive, data-driven implementations of RADI (and ADI)-based generalizations of balanced truncation, covering Linear Quadratic Gaussian (LQG) balanced truncation, $\mathcal{H}_{\infty}$ balanced truncation, positive-real balanced truncation, bounded-real balanced truncation, self-weighted balanced truncation, and balanced stochastic truncation. We also discuss practical considerations for improving computational efficiency. Numerical experiments demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
@@ 1 Introduction
Model order reduction (MOR) refers to system-theoretic techniques used to obtain simplified models that retain the essential input-output behavior of high-dimensional dynamical systems. These reduced order models (ROMs) are constructed to preserve the key dynamics and properties of the original system while being much lower in dimension. As a result, they offer significant computational advantages, making simulation, control, and analysis more efficient. For a detailed overview of MOR methods, refer to [5, 6, 7, 8].
Among MOR techniques, balanced truncation (BT) is a widely adopted and effective approach for linear systems [9]. It ensures that the ROM preserves the stability of the original system and provides guaranteed a priori error bound. BT truncates states that contribute minimally to the system’s input-output behavior, as determined by the system’s Hankel singular values. This results in a ROM that closely approximates the full-order model.
The main computational burden in BT lies in solving high-order Lyapunov equations to obtain the system Gramians. Numerous efficient algorithms, as discussed in surveys [10, 11], have been proposed for this purpose. These approaches require access to the system’s full state-space model and are therefore considered intrusive. In contrast, non-intrusive methods do not rely on internal state-space information but instead use system response data, such as transfer function samples or impulse responses [12, 13, 14, 15]. A notable example is the Quadrature-BT (QuadBT) algorithm introduced in [1], which constructs the ROM using numerical integration based on transfer function evaluations along the $j\omega$-axis or on sampled impulse responses and their derivatives. Over the last twenty years, the low-rank Alternating Direction Implicit (ADI) method has become a powerful tool for reducing the computational cost of BT in large-scale settings [16]. Recently, a data-driven, non-intrusive implementation of the ADI-based BT, which constructs the ROM using transfer function samples taken from the right half of the s-plane, was proposed in [2].
@@
BT has been generalized to various classes of dynamical systems [17], including descriptor systems [18], parametric systems [19], nonlinear systems [20], and second-order systems [21], while maintaining important properties such as passivity [22], contractivity [23], and the minimum-phase condition [24]. Recently, QuadBT was extended to include certain generalizations of BT, such as balanced stochastic truncation (BST) [24], positive-real BT (PR-BT) [24], and bounded-real BT (BR-BT) [23, 25]. Although this extension [3] is theoretically non-intrusive, it relies on access to samples of spectral factorizations of the transfer function. As practical techniques for acquiring such samples are not yet available, the method remains primarily of theoretical interest.
In this work, we introduce non-intrusive, data-driven interpolation algorithms that not only interpolate the original transfer function at chosen interpolation points but also preserve key system properties such as positive-realness, bounded-realness, and the minimum-phase condition. Additionally, we present an ADI-based non-intrusive data-driven framework applicable to several BT generalizations, including Linear Quadratic Gaussian-BT (LQG-BT) [26], $\mathcal{H}_{\infty}$-BT [27], PR-BT [24], BR-BT [23], self-weighted BT (SW-BT) [28], and BST [24]. In contrast to the quadrature-based approach in [3], the proposed methods rely on transfer function samples that can be practically obtained using frequency-domain techniques from the system identification literature [29]. The data-driven algorithms for LQG-BT and $\mathcal{H}_{\infty}$-BT proposed in this paper provide a control design approach that can be used to design LQG and $\mathcal{H}_{\infty}$-infinity controllers for unknown plants using only samples of the plant’s transfer function. Numerical results using a benchmark test model demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
@@
The paper is structured as follows. Section 2 introduces the MOR problem and provides a brief overview of key MOR techniques relevant to this work. Section 3 presents the main contributions. In Subsection 3.1, we propose non-intrusive data-driven implementations of LQG-BT and $\mathcal{H}_{\infty}$-BT. Subsection 3.2 introduces a data-driven implementation of PR-BT. Subsection 3.3 focuses on a data-driven BR-BT implementation. In Subsection 3.4, we propose a data-driven implementation of SW-BT. Subsection 3.5 presents a data-driven implementation of BST. Section 4 addresses practical considerations to prevent numerical issues during implementation. Section 5 evaluates the performance of the proposed algorithms through a benchmark numerical example. Finally, Section 6 offers concluding remarks.
@@ 2 Preliminaries
Consider a stable $n^{th}$-order linear time-invariant (LTI) system $G(s)$, represented in minimal state-space form as
@@
with matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $C\in\mathbb{R}^{p\times n}$, and $D\in\mathbb{R}^{p\times m}$.
Consider an $r^{th}$-order stable reduced-order approximation $\hat{G}(s)$ of $G(s)$ with minimal state-space realization:
@@
where $\hat{A}\in\mathbb{C}^{r\times r}$, $\hat{B}\in\mathbb{C}^{r\times m}$, $\hat{C}\in\mathbb{C}^{p\times r}$, and $r\ll n$.
The ROM $\hat{G}(s)$ is constructed from $G(s)$ through Petrov-Galerkin projection:
$\hat{A}=\hat{W}^{*}A\hat{V},\quad\hat{B}=\hat{W}^{*}B,\quad\hat{C}=C\hat{V}$,
@@
2.1 Interpolation Theory [31]
Consider the right interpolation points $(\sigma_{1},\dots,\sigma_{v})$ and left interpolation points $(\mu_{1},\dots,\mu_{w})$. In the interpolation framework, the projection matrices $\hat{V}\in\mathbb{C}^{n\times vm}$ and $\hat{W}\in\mathbb{C}^{n\times wp}$ are constructed as:
@@ 2.1 Interpolation Theory [31]
To enforce the Petrov-Galerkin condition ($\hat{W}^{*}\hat{V}=I$), $\hat{W}$ is replaced with $\hat{W}T_{w}$, where $T_{w}=(\hat{V}^{*}\hat{W})^{-1}$. The resulting ROM satisfies the interpolation conditions:
$G(\sigma_{i})=G_{r}(\sigma_{i}),\quad G(\mu_{i})=G_{r}(\mu_{i}).$ (3)
Additionally, if there are common right and left interpolation points, i.e., $\sigma_{i}=\mu_{j}$, the following Hermite interpolation conditions are also satisfied for those points:
$G^{\prime}(\sigma_{i})=G_{r}^{\prime}(\sigma_{i}).$ (4)
2.2 Interpolatory Loewner framework [12]
The matrix $D$ can be obtained by sampling the transfer function $G(s)$ at a sufficiently high frequency, since $D=\lim_{s\to\infty}G(s)$. Let $H(s)$ be defined as:
@@ 2.2 Interpolatory Loewner framework [12]
Samples of $H(s_{i})$ can be computed from those of $G(s_{i})$ as:
@@
In the Loewner framework, the following matrices are constructed from $H(s)$ samples at interpolation points:
@@
$\hat{W}^{*}B$ $=\begin{bmatrix}H(\mu_{1})\\ \vdots\\ H(\mu_{w})\end{bmatrix},\quad C\hat{V}=\begin{bmatrix}H(\sigma_{1})&\cdots H(\sigma_{v})\end{bmatrix}$, (5)
where $\hat{V}$ and $\hat{W}$ are as defined in (1) and (2), respectively. When $\sigma_{j}\approx\mu_{i}$,
@@
Thus, if the sets of right and left interpolation points share common elements, derivative samples $H^{\prime}(\sigma_{j})$ are also needed to construct $\hat{W}^{*}\hat{V}$ and $\hat{W}^{*}A\hat{V}$. The matrices $\hat{W}^{*}\hat{V}$ and $\hat{W}^{*}A\hat{V}$ exhibit a special structure known as the Loewner matrix and shifted Loewner matrix, respectively, which gives the name “Interpolatory Loewner Framework”. The matrix $\hat{W}^{*}\hat{V}$ is not always invertible, preventing the computation of $\hat{A}=(\hat{W}^{*}\hat{V})^{-1}\hat{W}^{*}A\hat{V}$ and $\hat{B}=(\hat{W}^{*}\hat{V})^{-1}\hat{W}^{*}B$. Instead, the ROM is constructed in descriptor form:
@@ 2.3 Pseudo-optimal Rational Krylov Algorithm (PORK) [32]
The projection matrices $\hat{V}$ and $\hat{W}$ in (1) and (2) solve:
@@
Pre-multiplying (7) with $\hat{W}^{*}$ under the condition $\hat{W}^{*}\hat{V}=I$ reveals that $\hat{A}=S_{v}-\hat{B}L_{v}$. This parameterizes $A_{r}$ in terms of $\hat{B}=\zeta$ while preserving the interpolation conditions induced by $\hat{V}$, i.e., varying $\zeta$ is equivalent to varying $\hat{W}$.
Let us assume that all interpolation points $(\sigma_{1},\cdots,\sigma_{v})$ lie in the right half of the $s$-plane. Consider the observable pair $(S_{v},L_{v})$ that satisfies the Lyapunov equation:
@@
With the free parameter $\zeta=Q_{v}^{-1}L_{v}^{T}$, we obtain $\hat{A}=S_{v}-\hat{B}L_{v}=-Q_{v}^{-1}S_{v}^{*}Q_{v}$. The resulting ROM
@@
Similarly, pre-multiplying (8) with $\hat{V}^{*}$ shows that $\hat{A}$ can also be expressed as $\hat{A}=S_{w}-L_{w}\hat{C}$. This parameterizes $\hat{A}$ in terms of $\hat{C}=\zeta$ while preserving the interpolation conditions induced by $\hat{W}$, since changing $\zeta$ is equivalent to varying $\hat{V}$.
Let us assume that all interpolation points $(\mu_{1},\cdots,\mu_{w})$ lie in the right half of the $s$-plane. For the controllable pair $(S_{w},L_{w})$ that satisfies the Lyapunov equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}=0,$ (10)
choosing the free parameter $\zeta=L_{w}^{T}P_{w}^{-1}$ leads to the expression $\hat{A}=S_{w}-L_{w}\hat{C}=-P_{w}S_{w}^{*}P_{w}^{-1}$. The resulting ROM
@@
satisfies the $\mathcal{H}_{2}$-optimality condition given by $\frac{\partial}{\partial\hat{B}}\big(\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{2}}^{2}\big)=0$. This approach is referred to as Output PORK (O-PORK) throughout this paper.
2.4 Balanced Truncation (BT) Family
The controllability Gramian $P$ and observability Gramian $Q$ are defined through the frequency-domain integrals:

### v3

@@ Abstract
Quadrature-based approximation of Gramians in standard balanced truncation yields a non-intrusive, data-driven implementation that requires only transfer function samples on the imaginary axis, which can be measured experimentally. This idea has recently been extended to several generalizations of balanced truncation, including positive-real balanced truncation, bounded-real balanced truncation, and balanced stochastic truncation. However, these extensions require samples of some spectral factorizations on the imaginary axis, and no practical method exists to obtain such data experimentally. As a result, these non-intrusive implementations are mainly of theoretical interest at present.
This paper shows that if the Gramians in these generalizations are approximated via rational interpolation rather than numerical integration, the resulting non-intrusive implementations do not require spectral factorization samples. Instead, they rely only on transfer function samples. Based on this idea, non-intrusive implementations are first developed for several variants of balanced truncation, wherein the Gramians are approximated implicitly using low-rank Alternating Direction Implicit (ADI) methods for Lyapunov and Riccati equations. These formulations include Linear Quadratic Gaussian (LQG) balanced truncation, $\mathcal{H}_{\infty}$ balanced truncation, positive-real balanced truncation, bounded-real balanced truncation, self-weighted balanced truncation, and balanced stochastic truncation. They require transfer function samples in the right half of the $s$-plane, which cannot be measured experimentally.
Next, building on these results, novel data-driven non-intrusive implementations are proposed that require only transfer function samples on the imaginary axis. Hence, unlike the quadrature-based and ADI-based approaches, these non-intrusive formulations can be implemented using practically measurable data. Numerical results are presented for benchmark problems in model order reduction, which show that the proposed non-intrusive implementations achieve accuracy comparable to their intrusive counterparts.
@@ 1 Introduction
Model order reduction (MOR) addresses the problem of constructing a low-dimensional dynamical system that approximates an original high-dimensional system. The aim is to create a reduced-order model (ROM) whose input-output response approximates that of the full-order system within an acceptable error tolerance. This provides a systematic approach for generating computationally efficient surrogate systems to accelerate simulation, analysis, and control design for large-scale models. For comprehensive reviews of MOR techniques, see [52, 46, 37, 7].
Balanced truncation (BT) is a classical and effective MOR technique for linear time-invariant (LTI) systems [34]. It constructs a ROM by truncating states associated with small Hankel singular values, which correspond to weakly controllable and observable dynamics. The resulting ROM preserves the stability of the full-order model (FOM) and provides an a priori error bound.
The primary computational cost of BT stems from solving Lyapunov equations to compute the controllability and observability Gramians. Efficient low-rank solvers for this task are reviewed in [9, 51]. These methods require explicit access to the system’s state-space matrices and are therefore classified as intrusive. In contrast, non-intrusive approaches construct ROMs solely from input–output data, such as samples of the transfer function or impulse response [21, 31, 36, 19, 11, 43, 50, 18, 1]. One such method is Quadrature-based BT (QuadBT) [16], which constructs the ROM using numerical quadrature applied to transfer function samples along the imaginary axis ($j\omega$) or to sampled impulse response and their derivatives. QuadBT has been extended to several classes of dynamical systems, such as second-order systems [49, 59] and quadratic output systems [42]. Other non-intrusive approximate balancing approaches include [30, 56, 45, 39, 17, 4].
Over the past two decades, the low-rank Alternating Direction Implicit (ADI) method has emerged as a numerically efficient and effective solver for the Lyapunov equations in BT [8]. Recently, a non-intrusive implementation of ADI-based BT was proposed [61]. This approach constructs the ROM directly from transfer function samples taken in the right half of the $s$-plane.
@@
BT has been extended to various system classes, including descriptor systems [35], parametric systems [58], nonlinear systems [29], and second-order systems [48]. These extensions preserve various system properties such as passivity [40], contractivity [38], and the minimum-phase condition [20]. Recently, QuadBT has been extended to include certain generalizations of BT [47], such as balanced stochastic truncation (BST) [20], positive-real BT (PRBT) [20], and bounded-real BT (BRBT) [38, 15]. Although this generalization [47] is formulated in a non-intrusive setting, it depends on samples of spectral factorizations of the transfer function. Since practical methods for obtaining such samples in real-world settings are currently unavailable, the method remains largely theoretical.
In this work, we first introduce a low-rank ADI-based non-intrusive framework applicable to several BT generalizations, including Linear Quadratic Gaussian BT (LQGBT) [26], $\mathcal{H}_{\infty}$BT [33], PRBT [20], BRBT [38], self-weighted BT (SWBT) [60], and BST [20]. In contrast to the quadrature-based approach in [47], the ADI-based non-intrusive implementations rely on transfer function samples in the right half of the $s$-plane. Although such samples cannot be measured in an experimental setting, the results developed for the ADI-based implementations are later generalized to obtain projection-based non-intrusive implementations that rely on transfer function samples on the imaginary axis, which can be measured experimentally. This leads to data-driven implementations of approximate BT, LQGBT, $\mathcal{H}_{\infty}$BT, PRBT, BRBT, SWBT, and BST. Numerical results show that both non-intrusive formulations perform comparably to their intrusive counterparts.
@@
The paper is structured as follows. Section 2 introduces the MOR problem and provides a brief overview of key MOR techniques relevant to this work. Sections 3 and 4 present the main contributions. In Section 3, we propose non-intrusive low-rank ADI-based implementations of LQGBT, $\mathcal{H}_{\infty}$ BT, PRBT, BRBT, SWBT, and BST. Section 4 presents projection-based data-driven counterparts of the algorithms in Section 3. The numerical performance of the proposed non-intrusive algorithms is tested in Section 5. Finally, Section 6 offers concluding remarks.
@@ 2 Preliminaries
Consider a stable LTI system $G(s)$ of order $n$, given in minimal state-space form by
@@
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $C\in\mathbb{R}^{p\times n}$, and $D\in\mathbb{R}^{p\times m}$.
Let $\hat{G}(s)$ denote a stable ROM of order $r\ll n$ that approximates $G(s)$, with minimal realization
@@
where $\hat{A}\in\mathbb{C}^{r\times r}$, $\hat{B}\in\mathbb{C}^{r\times m}$, and $\hat{C}\in\mathbb{C}^{p\times r}$.
The ROM $\hat{G}(s)$ is obtained via a Petrov–Galerkin projection of the FOM $G(s)$ as:
$\hat{A}=\hat{W}^{*}A\hat{V},\qquad\hat{B}=\hat{W}^{*}B,\qquad\hat{C}=C\hat{V}$,
@@
2.1 Krylov subspace-based Rational Interpolation [6]
Let $(\sigma_{1},\dots,\sigma_{v})$ denote a set of right interpolation points and $(\mu_{1},\dots,\mu_{w})$ a set of left interpolation points. In the rational interpolation framework, the projection matrices $\hat{V}\in\mathbb{C}^{n\times vm}$ and $\hat{W}\in\mathbb{C}^{n\times wp}$ are defined as
@@ 2.1 Krylov subspace-based Rational Interpolation [6]
To satisfy the bi-orthogonality condition $\hat{W}^{*}\hat{V}=I$, the left basis is modified as $\hat{W}\leftarrow\hat{W}T_{w}$, with $T_{w}=(\hat{V}^{*}\hat{W})^{-1}$. The resulting ROM interpolates the FOM at the prescribed interpolation points:
$G(\sigma_{i})=\hat{G}(\sigma_{i}),\quad G(\mu_{i})=\hat{G}(\mu_{i}).$ (3)
Moreover, at any interpolation point that appears in both sets—i.e., when $\sigma_{i}=\mu_{j}$—the ROM also matches the first derivative of the FOM at that point, satisfying the Hermite interpolation condition:
$G^{\prime}(\sigma_{i})=\hat{G}^{\prime}(\sigma_{i}).$ (4)
2.2 The Loewner framework [31]
The feedthrough matrix $D$ can be determined by sampling the transfer function $G(s)$ at a sufficiently large frequency, as it holds that $D=\lim_{s\to\infty}G(s)$. Define the strictly proper part of the FOM as
@@ 2.2 The Loewner framework [31]
Given samples of $G(s)$ at selected points $s_{i}$, the corresponding samples of $H(s_{i})$ are obtained via
@@
In the Loewner framework, the following matrices are assembled using samples of $H(s)$ at the right interpolation points $\{\sigma_{j}\}_{j=1}^{v}$ and left interpolation points $\{\mu_{i}\}_{i=1}^{w}$:
@@
$\hat{W}^{*}B$ $=\begin{bmatrix}H(\mu_{1})\\ \vdots\\ H(\mu_{w})\end{bmatrix}$,
$C\hat{V}$ $=\begin{bmatrix}H(\sigma_{1})&\cdots H(\sigma_{v})\end{bmatrix}$, (5)
where $\hat{V}$ and $\hat{W}$ are the projection matrices defined in (1) and (2), respectively.
When a right and a left interpolation point are nearly identical (i.e., $\sigma_{j}\approx\mu_{i}$), the divided differences approach derivatives:
@@
Consequently, when the interpolation sets share common points, constructing $\hat{W}^{*}\hat{V}$ and $\hat{W}^{*}A\hat{V}$ also requires samples of the derivative $H^{\prime}(\sigma_{j})$. The matrices $\hat{W}^{*}\hat{V}$ and $\hat{W}^{*}A\hat{V}$ have a special structure and are referred to as the Loewner and shifted Loewner matrices, respectively.
The matrix $\hat{W}^{*}\hat{V}$ can be singular, in which case the standard reduced-order matrices $\hat{A}=(\hat{W}^{*}\hat{V})^{-1}\hat{W}^{*}A\hat{V}$ and $\hat{B}=(\hat{W}^{*}\hat{V})^{-1}\hat{W}^{*}B$ are not well defined. In such cases, the ROM is constructed in descriptor form as
@@ 2.3 Pseudo-optimal Rational Krylov Algorithm (PORK) [55]
Then the projection matrices $\hat{V}$ and $\hat{W}$, defined in (1) and (2), satisfy the following Sylvester equations:
@@
Pre-multiplying (7) by $\hat{W}^{*}$ and enforcing the bi-orthogonality condition $\hat{W}^{*}\hat{V}=I$ yields $\hat{A}=S_{v}-\hat{B}L_{v}$. This expresses the reduced-order state matrix $\hat{A}$ in terms of the free parameter $\hat{B}=\zeta$, which preserves the interpolation conditions enforced by $\hat{V}$. Therefore, varying $\zeta$ is equivalent to modifying $\hat{W}$.
Assume that all interpolation points $\sigma_{1},\dots,\sigma_{v}$ are located in the right half of the $s$-plane and that the pair $(S_{v},L_{v})$ satisfies the Lyapunov equation:
@@
If the pair $(S_{v},L_{v})$ is observable, setting the free parameter $\hat{B}=Q_{v}^{-1}L_{v}^{T}$ yields $\hat{A}=S_{v}-\hat{B}L_{v}=-Q_{v}^{-1}S_{v}^{*}Q_{v}$. The resulting ROM
@@
Similarly, pre-multiplying (8) by $\hat{V}^{*}$ reveals that $\hat{A}$ can be written as $\hat{A}=S_{w}-L_{w}\hat{C}$. This expresses $\hat{A}$ in terms of the free parameter $\hat{C}=\zeta$, which is equivalent to varying $\hat{V}$; the interpolation conditions induced by $\hat{W}$ are not affected.
Now, assume all interpolation points $\mu_{1},\dots,\mu_{w}$ lie in the right half of the $s$-plane and the pair $(S_{w},L_{w})$ satisfies the Lyapunov equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}=0.$ (10)
If $(S_{w},L_{w})$ is controllable, choosing the free parameter $\hat{C}=L_{w}^{T}P_{w}^{-1}$ yields $\hat{A}=S_{w}-L_{w}\hat{C}=-P_{w}S_{w}^{*}P_{w}^{-1}$. The resulting ROM
@@
satisfies the $\mathcal{H}_{2}$-optimality condition $\frac{\partial}{\partial\hat{B}}\bigl(\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{2}}^{2}\bigr)=0$. This approach is referred to as Output PORK (O-PORK) in the remainder of the paper.
2.4 Balanced Truncation (BT) and Its extensions
The controllability Gramian $P$ and observability Gramian $Q$ of the state-space realization $(A,B,C,D)$ can be expressed in the frequency domain via the integrals